OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-3, -4) and (0, 2)?

OpenStudy (anonymous):

(y - 8) = 2(x - 3) (y + 3) = -2(x + 4) (y - 2) = -2(x + 4) (y + 4) = 2(x + 3)

OpenStudy (anonymous):

first we need to find the slope by using the slope formula. slope(m) = (y2 - y1) / (x2 - x1) are you familiar with this formula....do you know how to use it to find the slope ?

OpenStudy (anonymous):

let me know when you want to participate